OpenStudy (anonymous):

A health food store sells oatmeal for $3.60 per pound and bran flakes for $4.80 per pound. How many pounds of each should be used to get a mixture of 30 pounds that sells for $4.00 a pound?

OpenStudy (anonymous):

rid plz explain this one

OpenStudy (anonymous):

3.6x+4.8y=4(30) This is the equation

OpenStudy (anonymous):

cost*amount+cost*amount=total cost* total amount

OpenStudy (anonymous):

Wait I didnt finish it

OpenStudy (anonymous):

Whats the solution?

OpenStudy (anonymous):

Nope i got the wrong equation

OpenStudy (anonymous):

i feel i got it right in the mixture there will be 20 pounds of Oat and 10 pounds of Barn

OpenStudy (anonymous):

3.6(30-x)+4.8(x)=4(30)

OpenStudy (anonymous):

thanks rid. b4 i had taken it into % so was wrong

OpenStudy (anonymous):

x=10

OpenStudy (anonymous):

3.6 x + 4.8 (1-x) = 4 x=2/3 So you need 2/3 of oatmeal and 1/3 of bran flakes

OpenStudy (anonymous):

10 lbs of oatmeal and 20 lbs of bran

OpenStudy (anonymous):

no rid it will be 20lbs of Oat and 10lbs of Bran

OpenStudy (anonymous):

cross check it ! @RID

OpenStudy (anonymous):

Hey I like the way Fernando figured it out. Cool

OpenStudy (anonymous):

Ya sheg you r correct. I always make these mistakes!

OpenStudy (anonymous):

yeah it was nice @ fernando

OpenStudy (anonymous):

Thanks Fernando